# Teuken-7B-Base & Teuken-7B-Instruct: Towards European LLMs

### v2

@@ A.5 Evaluation
Ours (Instruct) .566 .606 .654 .436 .566
Ours (Base) .509 .567 .640 .343 .483
@@
A.5.1 Toxicity
We evaluate our models compared to other instruction-tuned models on the PolygloToxicityPrompts benchmark Jain et al. (2024). We report results from $\textsc{PTP}_{\text{SMALL}}$. Our evaluation settings are the same as those detailed in Section 3.4 of Jain et al. (2024), except the repetition parameter $K$ which is set to $K=1$ in our setup. We further report toxicity and profanity metrics provided by the Perspective API Lees et al. (2022). For both attributes $i\in\{\text{Profanity},\text{Toxicity}\}$ we report the Empirical Probability denoted $\text{EP}_{i}$ as well as the Average denoted $\text{A}_{i}$ as defined in Section 3.4 of Jain et al. (2024). Tables 34 to 39 show the results by language as well as aggregated, for the languages German, English, French, Italian and Spanish.
Meta-Llama-3.1-8B-Instruct .174 .196 .173 .219
Mistral-7B-Instruct-v0.3 .139 .160 .140 .183
Salamandra-7B-Instruct .206 .216 .196 .230
Aya-23-8B .177 .197 .176 .218
Occiglot-7B-eu5-Instruct .202 .212 .193 .226
Bloomz-7B1 .100 .115 .106 .134
Ours (Instruct) .216 .223 .201 .230
@@ A.5.1 Toxicity
A.6 Instruction Tuning
In the following, we provide more insights regarding our instruction tuned model. Particularly, we give an overview of the training datasets that we employed for instruction tuning the model (Section A.6.1), and present additional evaluation results (Section A.6.2).
A.6.1 Instruction Tuning Datasets
Table 40 provides an overview of the employed instruction tuning datasets.
Subset Language Count Ø Turns
anon8231489123/ShareGPT_Vicuna_unfiltered EN 37.6K 4.6
MBZUAI/Bactrian-X EN 26.9K 1.0
Open-Orca/OpenOrca EN 26.9K 1.0
WizardLM/WizardLM_evol_instruct_70k EN 26.9K 1.0
WizardLM/WizardLM_evol_instruct_V2_196k EN 26.8K 1.0
sahil2801/CodeAlpaca-20k EN 12.1K 1.0
lmsys/lmsys-chat-1m EN 11.2K 1.5
HuggingFaceH4/ultrachat_200k EN 7.0K 3.2
MBZUAI/Bactrian-X DE 63.7K 1.0
FreedomIntelligence/evol-instruct-deutsch DE 55.9K 1.0
FreedomIntelligence/alpaca-gpt4-deutsch DE 47.5K 1.0
FreedomIntelligence/sharegpt-deutsch DE 5.8K 2.9
LeoLM/German_Songs DE 943 1.5
bjoernp/ultrachat_de DE 909 1.0
LeoLM/German_Poems DE 378 1.0
Table 40: Instruction-tuning training dataset composition.
A.6.2 Instruction Tuning Evaluation Results
Subset Language Count Ø Turns
CohereForAI/aya_collection Multilingual 39.9K 1.0
MBZUAI/Bactrian-X Multilingual 32.3K 1.0
CohereForAI/aya_dataset Multilingual 13.1K 1.0
anon8231489123/ShareGPT_Vicuna_unfiltered EN 37.7K 4.6
sahil2801/CodeAlpaca-20k EN 12.1K 1.0
lmsys/lmsys-chat-1m EN 11.2K 1.5
WizardLM/WizardLM_evol_instruct_V2_196k EN 8.1K 1.0
Open-Orca/OpenOrca EN 8.1K 1.0
WizardLM/WizardLM_evol_instruct_70k EN 8.1K 1.0
HuggingFaceH4/ultrachat_200k EN 6.9K 3.1
anon8231489123/ShareGPT_Vicuna_unfiltered transl. DE 37.6K 4.6
lmsys/lmsys-chat-1m transl. DE 11.2K 1.5
WizardLM/WizardLM_evol_instruct_70k transl. DE 8.1K 1.0
Open-Orca/OpenOrca transl. DE 8.1K 1.0
WizardLM/WizardLM_evol_instruct_V2_196k transl. DE 8.1K 1.0
HuggingFaceH4/ultrachat_200k transl. DE 6.9K 3.1
FreedomIntelligence/sharegpt-deutsch DE 5.7K 2.9
bjoernp/ultrachat_de DE 905 1.0
Table 41: Instruction-tuning training dataset composition for 24 EU languages.
In Figure 14, we present the evaluation results of MT-Bench-X Weber et al. (2024). For the discussion of the results, we refer to Section 7.6.
@@ A.6.2 Instruction Tuning Evaluation Results
A.7 Software
We selected our training framework based on efficiency in terms of throughput (TFLOP/s) and maintenance. Therefore, we decided to train our models based on a fork of Megatron-LM Korthikanti et al. (2023) that supports various scalability features ensuring efficient training of transformer-based decoder-only models. Implementation contributions include a tensor-model-parallel SwiGLU layer Ramachandran et al. (2017); Shazeer (2020); Raffel et al. (2020); Google (2020) and fused RMSNorm Zhang and Sennrich (2019) integration. In particular, we made use of 3D parallelism, i.e., data, tensor, and pipeline parallelism. Additionally, we used ZeRO Korthikanti et al. (2023) to reduce memory requirements further by sharding the optimizer state.
A.8 Training Infrastructure
We trained our models on a supercomputer, which comprises 936 compute nodes, each containing 4$\times$ NVIDIA A100 (40 GB) GPUs connected via NVLink3 intra-node and through Mellanox HDR200 InfiniBand (IB) inter-node.

### v3

@@ A.4 Evaluation
Teuken-7B-Instruct (Ours) .579 .586 .677 .463 .591
Teuken-7B-Base (Ours) .527 .552 .627 .458 .473
@@
A.4.1 Toxicity
We evaluate our models compared to other instruction-tuned models on the PolygloToxicityPrompts benchmark [17]. We report results from $\textsc{PTP}_{\text{SMALL}}$. Our evaluation settings are the same as those detailed in Section 3.4 of [17], except the repetition parameter $K$ which is set to $K=1$ in our setup. We further report toxicity and profanity metrics provided by the Perspective API [6]. For both attributes $i\in\{\text{Profanity},\text{Toxicity}\}$ we report the Empirical Probability denoted $\text{EP}_{i}$ as well as the Average denoted $\text{A}_{i}$ as defined in Section 3.4 of [17]. Tables 32 to 36 show the results by language as well as aggregated, for the languages German, English, French, Italian and Spanish.
Teuken’s low toxicity scores can be attributed to the combination of our extensive filtering process and instruction tuning. First, we removed harmful content using heuristic rules and perplexity scores from KenLM. The resulting model still indicated room for improvement in toxicity. Therefore, for the last part of the training data, consisting of two trillion tokens, we applied an additional machine learning-based filter specifically trained to detect adult content. Combined with subsequent instruction tuning, this significantly reduced Teuken’s overall toxicity.
@@ A.4.1 Toxicity
A.5 Instruction Tuning
In the following, we provide more insights regarding our instruction tuned model. Particularly, we give an overview of the training datasets that we employed for instruction tuning the model (Section˜A.5.2), and present additional evaluation results (Section˜A.5.3).
A.5.1 Training details - Hyperparameters
We display the utilized hyperparameters in Table˜37.
Hyper-Parameter SFT DPO
Epochs 2 1
Weight decay 0.1 0.0
Batch size 128 128
Warmup steps 4300 50
Learning rate 1e-5 3.25e-6
Learning rate schedule cosine cosine
Optimizer AdamW AdamW
Adam Beta1 0.9 0.9
Adam Beta2 0.95 0.95
Max. Sequence length 2048 1024
Beta (DPO) NA 0.1
Table 37: Hyperparameter configuration during post-training
A.5.2 Instruction Tuning Datasets
Dataset Sample Size Min Distance Languages
IFeval-like [63] 5000 0.0 EN
GSM8K [28] 8000 (EN) / 5000 (DE) 0.0 EN, DE
Winogrande [29] 4000 (per language) 0.0 EN, DE
ARC [41] 180 (per language) 0.0 EN, DE
HellaSwag [45] 2000 (per language) 0.0 EN, DE
Sigma 30000 (per language) 0.1 EN, DE
Sigma Evolved 30000 (per language) 0.2 EN, DE
Sigma Everyday conversations 2260 0.0 EN
OpenMath-instruct 2 [50] 20000 (EN) / 5000 (DE) 0.0 EN, DE
Teuken Guide 20000 (EN) / 5000 (DE) 0.0 EN, DE
Teuken Self Awareness SFT 1130 0.0 EN, DE
Table 38: SFT datasets including sample size, minimum distance, and used languages.
Dataset Sample Size Languages
SauerkrautLM-Fermented 3000 DE
dpo-mix-7k 7000 (EN) / 2500 (per other language) EN,DE,FR,IT
truthy-dpo 1016 (EN) / 500 (per other language) EN,DE,FR,IT
Winogrande 1500 EN
ARC 180 EN
HellaSwag 750 EN
Table 39: DPO datasets including sample size and languages.
Table˜38 and Table˜39 provide an overview of the employed instruction tuning datasets and DPO datasets, respectively.
A.5.3 Instruction Tuning Evaluation Results
In Figure˜20, we present additional evaluation results of MT-Bench-X [1].
@@ A.5.3 Instruction Tuning Evaluation Results
A.6 Software
We selected our training framework based on efficiency in terms of throughput (TFLOP/s) and maintenance. Therefore, we decided to train our models based on a fork of Megatron-LM [55] that supports various scalability features ensuring efficient training of transformer-based decoder-only models. Implementation contributions include a tensor-model-parallel SwiGLU layer [42, 66, 13, 61] and fused RMSNorm [12] integration. In particular, we made use of 3D parallelism, i.e., data, tensor, and pipeline parallelism. Additionally, we used ZeRO [55] to reduce memory requirements further by sharding the optimizer state.
A.7 Training Infrastructure
[…]